OpenStudy (anonymous):

Solve the system of linear equations below x + y = 4 2x + 3y = 0

OpenStudy (anonymous):

First solve for \(x\) in terms of one of the other variables. The first one is easier to do this. \[x = 4 - y\] Then you can plug that into the second: \[2(4 - y) + 3y = 0\] After you've found y, you can substitute that back in and find x.

OpenStudy (campbell_st):

multiply the 1st equation by 2 then subtract the 2nd equation, this will eliminate x and allow you to solve for y 2x + 2y = 8 - 2x + 3y = 0 ---------------

OpenStudy (anonymous):

so then you what would be the answer?o:

OpenStudy (campbell_st):

well what do you get when you do the subtraction... when you have a value for y, substitute it into either equation to find x